OpenStudy (anonymous):

Reggie is three less than four times his daughter's age. Reggie is 69; how old is his daughter?

OpenStudy (hoblos):

18

OpenStudy (anonymous):

hoblos thats not an option on my test

OpenStudy (anonymous):

ash2326 what did u get?

OpenStudy (ash2326):

let reggies's age be x and his daughter's age be y he is 3 less than 4 times his daughter's age 4 times his daughter age =4y three less than it=4y-3 this is reggie's age 4y-3=x x=69 4y-3=69 4y=72 y=18 his daughter's age is 18

OpenStudy (anonymous):

oh never mind hoblos i was looking at the wrong question thanks

OpenStudy (ash2326):

18*4-3=69

OpenStudy (hoblos):

hahaha dont worry hope you've got the way ;)